You have just purchased a new warehouse. To finance the purchase, you’ve arranged for a 30-year mortgage loan for 80 percent of the $2,700,000 purchase price. The monthly payment on this loan will be $13,400. What is the APR on this loan? The EAR? full length descriptions and answer please!
Answer #1

Information provided:

Purchase price= $2,700,000

Mortgage= 0.80*2,700,000= $2,160,00

Time= 30 years*12= 360 months

Monthly payment= $13,400

The interest rate of the loan is calculated by entering the below in a financial calculator:

PV= -2,160,000

N= 360

PMT= 13,400

Press the CPT key and I/Y to compute the interest rate.

The value obtained is 0.5268.

Therefore, the annual interest rate is 0.5268.%*12= 6.3214% 6.32%.

Effective annual rate is calculated using the below formula:

EAR= (1+i/n)^n-1

Where i is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.0632/12)^12-1

      = 1.0651-1

        = 0.0651*100= 6.51%

Therefore, the effective annual rate is 6.51%.

In case of any query, kindly comment on the solution